1.There are n molecules in a gas, the average number of molecules moving in a definite direction will be
2. Oxygen is 16 times heavier than hydrogen.If mean kinetic energy of oxygen molecules at a certain temperature is E, then the mean kinetic energy of hydrogen molecules at the same temperature will be
3.If pressure is kept constant, then at what temperature the root mean square velocity of hydrogen molecules will be double the value at NTP
(1)    819 K    (2)    1092 K
    (3)    273 K    (4)    546 K
21.    At the same temperature, the pressure of one gm mole hydrogen and argon will be
    (1)    less in hydrogen    (2)    same in both
    (3)    more in hydrogen    (4)    nothing can be said
27.    The ratio of root mean square velocities of the two types of molecules in a mixture of and gas will be
Nitrogen molecule is 14 times heavier than hydrogen molecule.The temperature at which the root mean square velocity of hydrogen molecules is equal to the rms velocity of nitrogen molecules will be
    (1)    27°C    (2)    21.4°C
    (3)    21.4 K    (4)    absolute zero
Which of the following average of velocities of ideal gas molecules will be non zero
    (1)        (2)    
    (3)        (4)    
Two gases each having pressure P, volume V and temperature T are mixed.If the mixture has volume V and temperature T, then its pressure will be
    (1)    2P    (2)    P
    (3)    P/2    (4)    4P
42.    For one gm-mole gas,PV/T will be equal to
    (1)    nearly 2 cal/K
    (2)    4.2 cal/K
    (3)    4.2 joule/K
    (4)    2 × 10^7 erg/K
46.    An ideal gas Cv=(3R/2)  is kept in a closed vessel of volume 83 ×10^–4m3 at a pressure of1.6 × 10^–4m3 and at a temperature of 600 K.If 2.49 × 10^4 joule heat is supplied to the vessel then the final temperature will be
    (1)    600 K    (2)    625 K
    (3)    650 K    (4)    675 K
52.    A container has a mixture of hydrogen and oxygen in the ratio of 1:5.The ratio of the mean kinetic energies of the hydrogen and oxygen molecules is
    (1)    1 : 16    (2)    1 : 4
    (3)    1 : 5    (4)    1 : 1
53.    On mixing one gm-mole of helium at 300 K with three gm-mole of oxygen of 400 K at constant pressure if there is no exchange of energy with atmosphere, the final temperature will be
    (1)    350 K    (2)    383 K
    (3)    361 K    (4)    325 K
55.    The average momentum of molecules of a gas at a temperature T will be
    A vessel of volume 5000 cc contains (1/20) mole of molecular nitrogen at 1800 K.If 30% of the molecules are now dissociated the pressure inside the vessel in Pa will be
 A vessel contains equal number of hydrogen and oxygen molecules at one atmospheric pressure.The average number of collisions per sec. at the walls of the vessel by hydrogen molecules will be how many times the average number of collisions per sec. at the walls of the vessel by oxygen molecules
    (1)    16 times    (2)    4 times
    (3)1/16 times    (4)    1/4  times
For the escape of oxygen molecules from the earth's surface, its temperature should be
    (1)     4.11 × 10^3 K    (2) 2.09 × 10^4 K 
    (3)    1.6 × 10^5 K    (4) 3 × 10^5 K
If the density of a gas at NTP is 0.178 gm/litre and its molecular weight is 4, then the kinetic energy per gm-mole of gas at NTP will be
    (1) 3.4 × 10^2 J        (2)3.4× 10^2 erg
    (3) 3.4 × 10^3 erg            (4) 3.4 × 10^3 J
A container is filled with one gm-mole oxygen at a pressure of one at m and temperature 27°C.It is assumed that molecules of the gas are moving with velocity VrmsThen Vrmswill be (1 Atoms = 105 N/m2 and k = 1.38 x 10-23 J/K)
    (1)4.8 × 102 m/s        (2)48 × 102 m/s
    (3)4.8 × 102 cm/s        (4)48 × 102 cm/s
In the above question, number of collisions persecond at the walls of area 1 m2 ofthe container by the molecules is 
    (1)1.9 × 10^25    (2) 1.9 × 10^26
    (3) 1.9 × 10^27    (4) 1.9× 10^28

Grade:12th pass

1 Answers

User
5 years ago
46.PV =nRT
P=8.3×10-3
V= 1.6×10 and T=300K
Put all these value you will get, n=16/3
Since volume of the container is not changing all the heat given is used in increasing its internal energy,
Q= nCv∆T
2.5×104 = 16/3×3/2×25/3×∆T
∆T=375
T2 - 300=375
T2 = 675K